Q1- A lottery wheel contains 15% winning tickets. Tom buys one ticket and later another one. what are the chances of:-

a) getting two winning tickets.

b) getting a win ticket the second time only.

c) Not win at all .

d) Winning at least once.
Answer #1

Q1) Let the probability of winning be p. Then p=0.15 (Given) and probability of losing q= 1-p = 0.85. Here the buying of tickets is independent. Hence we can use the result P(A\bigcap B) = P(A) * P(B)

(a) Probability of two winning tickets = P(First ticket wins)* P(Second ticket wins) = 0.15*0.15 =0.0225

(b) Probability of a win ticket the second only = P(first ticket loses) * P(second ticket wins) = 0.85*0.15 = 0.1275

(c) Probability of no win = P(first ticket loses) * P(second ticket loses) = 0.85*0.85 = 0.7225

(d) Probability of winning at least once = 1 - probability of no win = 1-0.7225 = 0.2775